10gm of ice cubes at 0°C are released in a tumbler (water     equivalent 55 g) at 40°C. Assuming that negligible heat is taken from the surroundings, the temperature of water in the tumbler becomes nearly (L = 80 cal/g)


        (a)    31°C


(b)    22°C


        (c)    19°C


(d)    15°C


Grade:11

1 Answers

Amit Saxena
10 years ago

(b)

Let the final temperature be T

Heat gained by ice = mL + m × s × (T- 0)

= 10 × 80 + 10 × 1 × T

Heat lost by water = 55 × 1 × (40 - T)

By using law of calorimetery,

800 + 10T = 55 × (40-T)

Þ T = 21.54° C = 22°C